A team of runners is needed to run 1/4-mile relay race.If each runner must run 1/16 of a mile,how many team members are needed?
Answer by stanbon(75887)

A team of runners is needed to run 1/4-mile relay race.If each runner must run 1/16 of a mile,how many team members are needed?
------
Ans:: (1/4)/(1/16) = 16/4 = 4 needed
------------
Cheers,
Stan H.
-----------